# Chaos and Synchronization in Financial Leverages Dynamics: Modeling Systemic Risk with Coupled Unimodal Maps

### v1

@@ III.1 The Model
$\phi_{t}=\frac{\sum_{i=1}^{N}(\lambda_{i,t}\pi_{i}-1)}{\gamma}$
which now depends on a convex combination of the leverages of the $N$ banks $\lambda_{1},\lambda_{2},\dots,\lambda_{N}$ in a mean field fashion. It now remains to provide an expression for the variance of the aggregate return of the asset $\sigma_{e,t}$. A reasonable hypothesis is to assume that the estimate made at time $t$ is a weighted average ($\omega\in[0,1]$ being the weight) of the previously made estimate of the same kind and a statistical estimate of aggregate returns made by observing the returns between t1t-1 and $t$. I.e.:
$\sigma_{e,t}^{2}=\omega\sigma_{e,t-1}^{2}+(1-\omega)\hat{\sigma}^{2}_{e,t}$ (3)
@@
$\hat{\sigma}^{2}_{e,t}$ $=\widehat{\text{Var}}\left[\sum_{k=1}^{\mathbb{n}}r_{t-1+\frac{k}{\mathbb{n}}}\right]$
which is the aggregated variance of the AR(1) taking place between t1t-1 and $t$ as a function of the estimates of the parameters $\phi_{t-1}$ and $\sigma_{\epsilon}^{2}$, namely $\hat{\phi}_{t-1}$ and $\hat{\sigma}^{2}_{\epsilon}$. In the $\mathbb{n}\rightarrow\infty$ limit, one has $\hat{\sigma}_{e,t}^{2}\approx\frac{\mathbb{n}\hat{\sigma}^{2}_{\epsilon}}{(1-\hat{\phi}_{t-1})^{2}}$, having introduced the ML estimates for $\sigma_{\varepsilon}$ and $\phi_{t-1}$. Moreover, in the $\mathbb{n}\rightarrow\infty$ limit one expects the limit $\lim_{\mathbb{n}\rightarrow\infty}\mathbb{n}\sigma_{\epsilon}^{2}\doteq\Sigma_{\epsilon}$ to exists. Indeed, as already noticed one may consider the AR(1) as the discretization (with discretization step $\frac{1}{\mathbb{n}}$) of an Orstein-Uhlenbeck process, and so by a scaling argument the existence of the above limit may be deduced. Combining 2 and 3 one thus has, in the $\mathbb{n}\rightarrow\infty$ limit, (for $i\in\{1,2,\dots,N\}$)
$\lambda_{i,t}=\left(\frac{\omega}{\lambda_{i,t-1}^{2}}+\frac{(1-\omega)\alpha^{2}\Sigma_{\epsilon}}{(1-\hat{\phi}_{t-1})^{2}}\right)^{-\frac{1}{2}}$ (4)
Further, for $\mathbb{n}$ large enough $\hat{\phi}_{t-1}$ is a Gaussian with mean $\phi_{t-1}$ and variance $\frac{1-\phi_{t-1}^{2}}{\mathbb{n}}$. Writing thus $\hat{\phi}_{t-1}=\phi_{t-1}+\eta_{t-1}$ where $\eta_{t-1}\sim\mathcal{N}\left(0,\frac{1-\phi_{t-1}^{2}}{\mathbb{n}}\right)$ and expanding Eq. 4 at zeroth order, it is possible to write, for ($i\in\{1,2,\dots,N\}$) for $i=1,2,\dots,N$:
$\lambda_{i,t}$ $=\left(\frac{\omega_{i}}{\lambda_{i,t-1}^{2}}+\frac{(1-\omega_{i})\alpha^{2}\Sigma_{\epsilon}}{\left(1-\phi_{t-1}\right)^{2}}\right)^{-\frac{1}{2}}\doteq T_{i}(\{\lambda_{i,t-1}\}_{i=1,2,\dots,N})$ (5)
which is the evolution law already introduced in the previous section. Note that, should one have considered the first order term in the expansion above, one would have obtained a heteroschedastic noise term superimposed on the deterministic skeleton given by the maps $T$. This kind of system has been studied in lillo2023analysis in case $N=1$. Here, dealing with the more general case in which $N\geq 1$, we aim to first understand the behavior of the deterministic skeleton before considering the random system that is based on it.
@@ III.2 Homogeneous case: sychronization
$\lim_{n\rightarrow\infty}|T_{1}^{n}(\lambda_{1,0},\lambda_{2,0})-T_{2}^{n}(\lambda_{1,0},\lambda_{2,0})|=0.$
@@ III.3 Heterogeneous case
The box counting dimension for the attractor shown above has been calculated to be equal to $1.203\pm 0.006$, thus confirming its apparently fractal nature (see the supplementary material for details). Finally, the nature of the attractor led us to consider applying the main theorem in wang2008toward about the existence of an SRB measure supported on it. However, it seems that not all of the hypotheses can be met in our case: The main issue is that for $|\omega_{1}-\omega_{2}|$ going to zero, although the attractor becomes essentially one dimensional, there is no collapse of dimensionality for $\omega_{1}=\omega_{2}$ and initial conditions approach the diagonal only asymptotically; whereas, for the Hénon map, when the parameter usually denoted with “bb” is zero there is a collapse of the dimensionality and all initial conditions are mapped to a 1D segment.
@@
Let’s focus here on the case with $\pi=0$. This case may be looked at as a forcing-forced system, in which the larger bank “drives” the smaller one. Alternatively, one may say that the dynamics of the smaller one is described by a random process where the randomness is generated by a deterministic unimodal map. This setting is indeed a subcase of the general “heterogeneous case” previously introduced, but the fact that several interesting results may be, even rigorously, obtained makes it an interesting case to study.
@@
$\lambda_{1,t+1}=f_{\lambda_{2,t}}(\lambda_{1,t})$
@@
$f_{y}(x)=\left(\frac{\omega_{1}}{x^{2}}+\frac{(1-\omega_{1})\gamma^{2}\alpha^{2}\Sigma_{\epsilon}}{(1+\gamma-y)^{2}}\right)^{-\frac{1}{2}}$
@@
This system is in fact a skew product system
$\lambda_{1,t+1}$ $=f_{\lambda_{2,t}(\lambda_{1,t})}$
$\lambda_{2,t+1}$ $=T(\lambda_{2,t}).$
To present some analytical results, it is useful to introduce the invertible extension of $T$, $\hat{T}:\hat{I}\rightarrow\hat{I}$ where
@@
$\hat{T}((\lambda_{2,i})_{i\in\mathbb{Z}})=(\lambda_{2,i+1})_{i\in\mathbb{Z}}.$
Also, for $\boldsymbol{\lambda}\in\hat{I}$, let’s write $f^{n}_{\boldsymbol{\lambda}}=f_{\lambda_{2,n-1}}\circ f_{\lambda_{2,n-2}}\circ\dots\circ f_{\lambda_{2,1}}\circ f_{\lambda_{2,0}}$. (Let us stress once more that the evolution of the forcing bank does not depend on the forced bank and has already been discussed when dealing with the case $N=1$.) Let’s now give some analytical results (refer to supplementary material for further details and proofs). To begin with, no matter the initial conditions, the leverages of the forced bank will always behave asymptotically in the same manner (determined by the initial condition of the large bank, as specified further in Theorem 6). This also means that if one has two or more banks with weights $\pi_{i}=0$, their orbits will synchronize under the common forcing of the bigger bank.
@@
(Synchronization on the Fiber) For any orbit $\boldsymbol{\lambda}\in\hat{I}$ of the forcing bank and for any initial conditions of the forced bank $\lambda_{1,0},\,\lambda_{1,0}^{\prime}\in[1,\infty)$
$\lim_{n\rightarrow\infty}|\lambda_{1,n}-\lambda_{1,n}^{\prime}|=0.$
@@
(Negative Lyapunov Exponent on the Fiber) For any $\boldsymbol{\lambda}\in\hat{I}$ and $\lambda_{1,0}\in[1,\infty)$
$\Lambda_{1}(\lambda_{1,0},\boldsymbol{\lambda})\doteq\lim_{n\rightarrow\infty}\frac{1}{n}\log{|(f^{n}_{\boldsymbol{\lambda}})^{\prime}(\lambda_{1,0})|}<0.$
@@
If instead the forcing is chaotic (meaning that it is topologically transitive on a finite union of intervals and that it has a unique a.c.i.p. with respect to Lebsegue), then the forced one shows some irregularity too:
@@
Next, it is possible to provide an explicit characterization of the behavior of the leverage of the small bank as a function of the past orbit of the forcing leverage. In particular, one has:
@@
Let $x(\boldsymbol{\lambda})$ be the random fixed point for bank 1, i.e. such that $f_{\lambda_{2,0}}(x(\boldsymbol{\lambda}))=x(\hat{T}(\boldsymbol{\lambda}))$. Then
$x(\boldsymbol{\lambda})=\frac{1}{\sqrt{\sum_{i=0}^{\infty}A(\lambda_{2,-1-i})\omega_{1}^{i}}}$

### v2

@@ III.1 The Model
$\phi_{t}=\frac{\sum_{i=1}^{N}(\lambda_{i,t}\pi_{i}-1)}{\gamma}$ (13)
which now depends on a convex combination of the leverages of the $N$ banks $\lambda_{1},\lambda_{2},\dots,\lambda_{N}$ in a mean field fashion. At time $t+1$, banks need to update their target leverage so that $\lambda_{i,t+1}=\frac{1}{\alpha\sigma_{e,t+1}}$. Thus, it now remains to provide an expression for the forecast of the variance of the aggregate return of the asset $\sigma_{e,t+1}$. A reasonable hypothesis is to assume that the forecast made by bank $i$ at time $t+1$ is a weighted average ($\omega_{i}\in[0,1]$ being the weight) of the previously made forecast of the same kind and a statistical estimate of aggregate returns made by observing the returns between $t$ and $t+1$. I.e., for bank $i$:
$\sigma_{e,t+1}^{2}=\omega_{i}\sigma_{e,t}^{2}+(1-\omega_{i})\hat{\sigma}^{2}_{e,t+1}$ (14)
and {align} ^σ^2_e,t+1=^Var[∑_k=1^nr_t+kn] which is the aggregated variance of the AR(1) taking place between $t$ and $t+1$ as a function of the estimates of the parameters $\phi_{t}$ and $\sigma_{\epsilon}^{2}$, namely $\hat{\phi}_{t}$ and $\hat{\sigma}^{2}_{\epsilon}$. In the $\mathbb{n}\rightarrow\infty$ limit, one has $\hat{\sigma}_{e,t+1}^{2}=\frac{\mathbb{n}\hat{\sigma}^{2}_{\epsilon}}{(1-\hat{\phi}_{t})^{2}}$ (see Appendix A of Mazzarisi et al. (2019) for the derivation of this result), having introduced the Maximum Likelihood estimates for $\sigma_{\varepsilon}$ and $\phi_{t}$. These are given by (see again Mazzarisi et al. (2019)):
$\hat{\phi}_{t}=\frac{\sum_{k=1}^{\mathbb{n}}r_{t+k/\mathbb{n}}r_{t+(k-1)/\mathbb{n}}}{\sum_{k=1}^{\mathbb{n}}r^{2}_{t+(k-1)/\mathbb{n}}}$ (15)
@@
$\hat{\sigma}_{\epsilon}^{2}=\frac{\sum_{k=1}^{\mathbb{n}}(r_{t+k/\mathbb{n}}-\hat{\phi}_{t}r_{t+(k-1)/\mathbb{n}})^{2}}{\mathbb{n}}$ (16)
Moreover, in the $\mathbb{n}\rightarrow\infty$ limit one expects the limit $\lim_{\mathbb{n}\rightarrow\infty}\mathbb{n}\sigma_{\epsilon}^{2}\doteq\Sigma_{\epsilon}$ to exists. Indeed, as already noticed one may consider the AR(1) as the discretization (with discretization step $\frac{1}{\mathbb{n}}$) of an Orstein-Uhlenbeck process, and so by a scaling argument the existence of the above limit may be deduced. Combining equations 6 and 14 one thus has, in the $\mathbb{n}\rightarrow\infty$ limit, (for $i\in\{1,2,\dots,N\}$)
$\lambda_{i,t+1}=\left(\frac{\omega_{i}}{\lambda_{i,t}^{2}}+\frac{(1-\omega_{i})\alpha^{2}\Sigma_{\epsilon}}{(1-\hat{\phi}_{t})^{2}}\right)^{-\frac{1}{2}}$ (17)
Further, for $\mathbb{n}$ large enough $\hat{\phi}_{t}$ is a Gaussian with mean $\phi_{t}$ and variance $\frac{1-\phi_{t}^{2}}{\mathbb{n}}$. Writing thus $\hat{\phi}_{t}=\phi_{t}+\eta_{t}$ where $\eta_{t}\sim\mathcal{N}\left(0,\frac{1-\phi_{t}^{2}}{\mathbb{n}}\right)$ and expanding Eq. 17 at zeroth order, it is possible to write, for $i=1,2,\dots,N$:
[…]
which is the evolution law already introduced in the previous section. Note that, should one have considered the first order term in the expansion above, one would have obtained a heteroschedastic noise term superimposed on the deterministic skeleton given by the maps $T$. This kind of system has been studied in Lillo et al. (2023a) in case $N=1$. Here, dealing with the more general case in which $N\geq 1$, we aim to first understand the behavior of the deterministic skeleton before considering the random system that is based on it.
@@ III.2 Homogeneous case: sychronization
$\lim_{n\rightarrow\infty}|T_{1}^{n}(\lambda_{1,0},\lambda_{2,0})-T_{2}^{n}(\lambda_{1,0},\lambda_{2,0})|=0.$ (19)
@@ III.3 Heterogeneous case
The box counting dimension for the attractor shown above has been calculated to be equal to $1.203\pm 0.006$, thus confirming its apparently fractal nature (see the supplementary material for details). Finally, the nature of the attractor led us to consider applying the main theorem in Wang and Young (2008) about the existence of an SRB measure supported on it. However, it seems that not all of the hypotheses can be met in our case: The main issue is that for $|\omega_{1}-\omega_{2}|$ going to zero, although the attractor becomes essentially one dimensional, there is no collapse of dimensionality for $\omega_{1}=\omega_{2}$ and initial conditions approach the diagonal only asymptotically; whereas, for the Hénon map, when the parameter usually denoted with “bb” is zero there is a collapse of the dimensionality and all initial conditions are mapped to a 1D segment.
@@
Let’s focus here on the case with $\pi=0$. This case may be looked at as a forcing-forced system, in which the larger bank “drives” the smaller one. This setting is indeed a subcase of the general “heterogeneous case” previously introduced, but the fact that several interesting results may be, even rigorously, obtained makes it an interesting case to study.
@@
$\lambda_{1,t+1}=f_{\lambda_{2,t}}(\lambda_{1,t})$ (20)
@@
$f_{y}(x)=\left(\frac{\omega_{1}}{x^{2}}+\frac{(1-\omega_{1})\gamma^{2}\alpha^{2}\Sigma_{\epsilon}}{(1+\gamma-y)^{2}}\right)^{-\frac{1}{2}}$ (21)
@@
This system is in fact a skew product system {align} λ_1,t+1=f_λ_2,t(λ_1,t)
λ_2,t+1=T(λ_2,t). To present some analytical results, it is useful to introduce the invertible extension of $T$, $\hat{T}:\hat{I}\rightarrow\hat{I}$ where
@@
$\hat{T}((\lambda_{2,i})_{i\in\mathbb{Z}})=(\lambda_{2,i+1})_{i\in\mathbb{Z}}.$ (23)
Also, for $\pmb{\lambda}\in\hat{I}$, let’s write $f^{n}_{\pmb{\lambda}}=f_{\lambda_{2,n-1}}\circ f_{\lambda_{2,n-2}}\circ\dots\circ f_{\lambda_{2,1}}\circ f_{\lambda_{2,0}}$. (Let us stress once more that the evolution of the forcing bank does not depend on the forced bank and has already been discussed when dealing with the case $N=1$.) Let’s now give some analytical results (refer to supplementary material for further details and proofs). To begin with, no matter the initial conditions, the leverages of the forced bank will always behave asymptotically in the same manner (determined by the initial condition of the large bank, as specified further in Theorem 6). This also means that if one has two or more banks with weights $\pi_{i}=0$, their orbits will synchronize under the common forcing of the bigger bank.
@@
(Synchronization on the Fiber) For any orbit $\pmb{\lambda}\in\hat{I}$ of the forcing bank and for any initial conditions of the forced bank $\lambda_{1,0},\,\lambda_{1,0}^{\prime}\in[1,\infty)$
$\lim_{n\rightarrow\infty}|\lambda_{1,n}-\lambda_{1,n}^{\prime}|=0.$ (24)
@@
(Negative Lyapunov Exponent on the Fiber) For any $\pmb{\lambda}\in\hat{I}$ and $\lambda_{1,0}\in[1,\infty)$
$\Lambda_{1}(\lambda_{1,0},\pmb{\lambda})\doteq\lim_{n\rightarrow\infty}\frac{1}{n}\log{|(f^{n}_{\pmb{\lambda}})^{\prime}(\lambda_{1,0})|}<0.$ (25)
@@
If instead the forcing is chaotic (meaning that it is topologically transitive on a finite union of intervals and that it has a unique a.c.i.m. with respect to Lebsegue), then the forced one shows some irregularity too:
@@
Next, it is possible to provide an explicit characterization of the behavior of the leverage of the small bank as a function of the past orbit of the forcing leverage. First, put, for $z\in I$:
$a_{z}\doteq\frac{1+\gamma-z}{\gamma\alpha\sqrt{(1-\omega_{1})\Sigma_{\epsilon}}}$ (26)
and define the function $x(\pmb{\lambda}):\hat{I}\rightarrow\mathbb{R}$
$x(\pmb{\lambda})\doteq\limsup_{n\rightarrow\infty}f^{n}_{\hat{T}^{-n}(\pmb{\lambda})}(a_{\lambda_{2},-n})$ (27)
Where $\pmb{\lambda}=(\lambda_{2,i})_{i\in\mathbb{Z}}\in\hat{I}$. Then one has:
@@
For every $\pmb{\lambda}\in\hat{I}$, it holds $f_{\lambda_{2,0}}(x(\pmb{\lambda}))=x(\hat{T}(\pmb{\lambda}))$ and
$x(\pmb{\lambda})=\frac{1}{\sqrt{\sum_{i=0}^{\infty}A(\lambda_{2,-1-i})\omega_{1}^{i}}}$ (28)
@@
Supplementary Material: Proofs of theorems, as well as additional results and numerical simulations, are provided in the supplementary material file.